Find the average velocity of the equation
x=t²+2t+3 when t=0, t=5
1
Expert's answer
2019-06-14T13:51:50-0400

By definition, the average velocity is just a ratio of a distance travelled to the time it takes to travel this distance:


"\lt v \gt =\frac{x(t_2)-x(t_1)}{t_2 - t_1}"

In the given system of units (no dimension is specified) we have:


"t_2 = 5, \quad t_1 = 0\\
x(t_2) = 25 + 10 + 3 = 38, \quad x(t_1) = 3"

Hence, we obtain:


"\lt v \gt =\frac{38-3}{5} = 7"

Answer: 7